# Adaptive Blind All-in-One Image Restoration

### v1

@@ Abstract
Blind all-in-one image restoration models aim to recover a high-quality image from an input degraded with unknown distortions. However, these models require all the possible degradation types to be defined during the training stage while showing limited generalization to unseen degradations, which limits their practical application in complex cases. In this paper, we propose a simple but effective adaptive blind all-in-one restoration (ABAIR) model, which can address multiple degradations, generalizes well to unseen degradations, and efficiently incorporate new degradations by training a small fraction of parameters. First, we train our baseline model on a large dataset of natural images with multiple synthetic degradations, augmented with a segmentation head to estimate per-pixel degradation types, resulting in a powerful backbone able to generalize to a wide range of degradations. Second, we adapt our baseline model to varying image restoration tasks using independent low-rank adapters. Third, we learn to adaptively combine adapters to versatile images via a flexible and lightweight degradation estimator. Our model is both powerful in handling specific distortions and flexible in adapting to complex tasks, it not only outperforms the state-of-the-art by a large margin on five- and three-task IR setups, but also shows improved generalization to unseen degradations and also composite distortions.
@@ 1 Introduction
Figure 1: Our model significantly outperforms state-of-the-art all-in-one image restoration (IR) methods, Restormer [63], PromptIR [40], and DiffUIR [66], across five known IR tasks, three unseen tasks, and three mixed degradation scenarios. The plot is normalized along each axis, with the lowest value positioned on the second circle and the highest value on the outermost circle.
@@
The inherently ill-posed nature of IR presents a significant challenge for conventional approaches limiting their effectiveness [38, 22, 34, 21, 47, 12, 18, 15]. Recent advances in deep learning techniques have led to remarkable progress in IR [27, 63, 40, 41, 11], achieving substantial improvements in reconstruction accuracy. However, these frameworks demonstrate superior performance only in modeling dedicated degradations, as they are typically trained on datasets with individual and specific IR tasks.
@@
To address the limitations of using separate models for each degradation type, recent image restoration (IR) models have adopted an all-in-one approach, i.e., they are designed to handle multiple degradation types within a single model, alleviating the dependency on dedicated models for various IR tasks. While these models have demonstrated success in integrating diverse IR tasks into a unified framework, several challenges persist when applied to real-world IR problems. First, these models assume that the type of distortions are known in advance, allowing them to specifically target and remove them [63]. In practice, however, IR must operate in a blind setting, with no prior knowledge of the degradation present in a given image, making blind IR significantly more challenging. Second, real images frequently contain composite degradations — e.g., a moving vehicle in a low-light scene, or a compressed image captured in hazy weather. Although existing models are designed to handle diverse degradation types, they typically process only a single type of distortion at a time, limiting their effectiveness in practical cases. Third, current all-in-one methods require access to all degradation types during training, which restricts their generalizability to unseen degradations. Adapting these models to new, unobserved degradation types while retaining the all-in-one functionality typically necessitates retraining the entire model with the expanded set of degradations — a process that is both computationally expensive and time-consuming.
@@
Contribution: This paper presents an Adaptive Blind All-in-One Image Restoration (ABAIR) method, designed to bridge the gap between IR techniques and their application in practical complex scenarios. ABAIR effectively addresses multiple and composite degradations with a flexible structure that can be easily updated to include new degradations. Our approach is a simple yet effective scheme that combines three main components. First, we propose a large-scale pre-training with synthetic degradations to obtain a robust weight initialization. We propose a modified CutMix [61] to generate images under multiple distortions with a segmentation head for degradation recognition. Second, we learn disentangled restoration modules by training independent adapters — specifically LoRA [19] — to bridge the gap between synthetic and real-world data. Third, to derive a blind all-in-one IR method, we learn a lightweight image-degradation estimator to select the most suitable combination of adapters based on the input image. With the pre-training on synthetic degradations and the blending operation, our model achieves the ability to handle composite distortions within a single image. Moreover, the adapter-based design enables a flexible model structure that can be easily updated to address new distortions. We can simply train a new adapter module for the added degradation and retrain the lightweight estimator — without losing prior knowledge on other IR tasks. Our approach offers a blind all-in-one model for versatile IR, achieving superior performance on all-in-one restoration benchmarks including three- and five- degradation settings, generalizing to three unseen degradations and handling mixed degradations, as shown in Fig. 1.
@@ 2 Related Work
Recently, multi-degradation and all-in-one IR approaches have gained significant attention. Multi-degradation methods [63, 5, 62, 60, 39] proposed unified model architectures that are effective on multiple IR tasks. However, these methods are trained in a way that one set of parameters can handle only one specific type of degradation. Therefore, one still needs to allocate different weights (though with the same architecture) for different degradations. Furthermore, to restore a versatile image, the type of degradation has to be known so that the corresponding parameters can be loaded, this non-blind scheme further hinders the efficiency and effectiveness for real-world applications.
@@
Meanwhile, blind all-in-one IR approaches utilize specialized modules to blindly distinguish degradation types. For instance, AirNet [27] uses a contrastive-based encoder to extract a latent degradation representation from the input images. X-Restormer [8] adds a spatial self-attention module to the transformer block to enhance spatial mapping capabilities. IDR [64] learns degradation-specific priors and incorporates them into restoration, while DiffUIR [66] equips a diffusion model with a condition guidance. Despite the ability to blindly process images, these methods are still incapable of dealing with composite distortions or unseen distortions, due to their focus on dedicated degradations and standard IR benchmarks.
@@
Figure 2: General schema of our proposed method. Our method is divided into three phases. In Phase I we pre-train our baseline model with synthetic degradations of high-fidelity images. Each image contains different degradations in different regions, and a segmentation head learns to predict them, while a restoration loss aims at restoring the image. In this way, the model is able to distinguish and generalize well to multiple degradations. In Phase II, we learn degradation-specific adaptors using standard image restoration datasets. In Phase III, we learn a lightweight degradation estimator to adaptively blend the adapters based on the degradation profile of the input image. This 3-phase methodology makes our method flexible to deal with images containing multiple distortions and easy to update for new ones as it only requires training an adapter for the new distortion and retraining the degradation estimator.
@@
Prompt learning techniques, which capture task-specific context, have shown promise in guiding adaptation for vision tasks [20]. Leveraging this concept, recent methods encode degradation-specific information and guide the restoration model to enhance low-fidelity images with more than one possible distortion. PromptIR [40], for instance, integrates a dedicated prompt block to capture degradation-specific features from input images, while DA-CLIP [30], MPerceiver [2], ProRes [31] and Painter [50] leverage large pre-trained models as prompt generators. The latter approaches, however, are often constrained by high memory demands of large models. Additionally, existing all-in-one approaches require all degradation types to be predefined during training, limiting the addition of new distortion to already trained models.
In contrast, aiming at versatile image restoration, our model is developed to blindly process images with various kinds of distortion, either in single or composite forms. With the mixed design of lightweight degradation adapters, our model is also flexible and extensible to unseen distortions, thus shows great generalizability towards the challenging blind all-in-one IR problem.
@@
Figure 3: Examples of our synthetic degradation generation for five traditional distortions.
@@
Fine-tuning models by updating all parameters is computationally inefficient, especially for large-scale models. Parameter Efficient Fine-Tuning (PEFT) addresses this by reducing the number of trainable parameters and memory usage while achieving comparable performance to full fine-tuning [54]. Low-Rank Adaptation (LoRA) [19] introduced a reparametrization strategy for fine-tuned weights, where the weights of a specific layer $W^{\prime}=W+\Delta W$ are represented as a linear combination of the pre-trained weights $W$ and the updated term $\Delta W$ obtained through a low-rank decomposition, $\Delta W=BA$. Here, $B\in\mathbb{R}^{d\times k}$, $A\in\mathbb{R}^{r\times k}$ with rank $r\ll\text{min}(d,k)$. By optimizing only the low-rank matrices $A$ and $B$, LoRA reduces the number of parameters required for adapting that layer (vs tuning all parameters), minimizing memory and computational demands. Building upon LoRA, alternative decomposition methods have been proposed, such as Vector-based Random Matrix Adaptation [24] and Conv-LoRA[67].
PEFT techniques have proven valuable not only for domain adaptation in large models but also for applications like task arithmetic and continual learning, mitigating the problem of catastrophic forgetting—where models lose previously acquired knowledge upon learning new tasks[10]. In IR, Park et al. [39] applied low-rank decomposition within single-task IR frameworks to enhance model performance. In contrast, we propose a simple but novel approach to merge different adaptors — LoRAs — to obtain a flexible blind all-in-one model capable of learning new tasks without forgetting previously learned IR tasks. This is achieved through a disentangled scheme that requires minimal retraining for each new task.
@@ 3 Method
Figure 2 presents an overview of our proposed approach for enhancing low-quality images by systematically addressing the key limitations in current IR methods. Our approach consists of three phases, roughly targeting the following challenges: (i) robust generalization to varying types of degradations (Phase I), (ii) effective adaptation to specific degradations (Phase II), (iii) flexible all-in-one mechanism able to handle unknown and mixed degradations (Phase III). Note that we freeze the parameters of baseline after Phase I and adapters after and II, thus preserving maximal knowledge learned from previous tasks and avoid catastrophic forgetting. This design also allows us to adapt to new distortions with minimal additional training as only a new adapter and the light-weighted estimator will be trained.
Phase I: Pre-training with synthetic degradations.
Recent advances demonstrate that large-scale pre-trained models can significantly improve performance across various tasks [1]. Moreover, pre-training with synthetic data has also shown effectiveness when the domain gap is properly addressed [55]. We hypothesize that applying large-scale pre-training to IR can yield a notable performance boost compared to training solely on traditional IR datasets. Thus, as depicted in Fig. 2, Phase I involves training a baseline IR model with large-scale data with synthetic degradations to provide a robust weight initialization for handling complex degradations.
To obtain large-scale data for pre-training, we define a degradation pipeline that generates a low-quality counterpart for each high-quality input image by introducing different synthetic distortions. We focus on five common distortions: noise, blur, rain, haze, and low-light conditions. Each synthetic distortion closely approximates its real-world counterpart. For instance, low-light conditions are simulated by compressing the image histogram, reducing the dynamic range of pixel intensities. In contrast, haze is introduced by adding an achromatic layer based on the depth map estimated from DepthAnythingv2 [55]. Figure 3 illustrates examples of each synthetic degradation. We use the Google Landmarks dataset (GLD) [53], a large-scale collection of five million images of diverse landmarks worldwide. To ensure that the inputs to our pipeline are high quality, we filter out the images with a resolution smaller than 400 pixels on any side and exclude images with a NIMA score [46] below the 4.90 threshold, retaining only high-quality natural images.
@@
While training on single distortions enables the model to address each type individually, real-world IR often involves mixed degradations. However, directly applying multiple distortions to an image often leads to severe quality degradation, making it difficult for the model to learn to reconstruct the original image. To tackle this, we implement two strategies: (i) a degradation CutMix [61] technique, and (ii) a segmentation head with a cross-entropy loss. Our first strategy — inspired by Yang et al. [55] for depth estimation — applies two different degradations divided into different regions. This setup helps the model distinguish and manage multiple degradation types within the same image. As our second strategy, we incorporate a segmentation head into the baseline model, which outputs a per-pixel map of the distortions. By comparing this map with the ground truth, we guide the model to recognize and differentiate multiple types of degradation within a single image. We report additional details about our baseline and the generation of the synthetic distortions in the Supplementary Material.
Table 1: 5-degradations setup. Quantitative results on five IR datasets comparing the state-of-the-art all-in-one methods and our approach. Ours (Oracle) is an upper bound for our approach: it computes the best reachable value in case our estimator always chooses the correct degradation.
@@
Ours OH 37.73 .978 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.15 .920 59M
Ours SW 37.79 .979 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.17 .921 59M
Ours (Oracle) 39.09 0.981 33.54 .984 31.40 0.901 29.10 .879 24.45 .866 31.39 .922 59M
@@
By generating distortions synthetically, we achieve a robust weight initialization for our baseline model. However, a domain gap persists between our pre-training data (due to both the different dataset and the synthetic distions), standard IR datasets, and real-world conditions. For instance, accurately simulating haze is particularly challenging due to its dependence on light scattering by particles, which varies with depth and atmospheric conditions. To bridge this gap, we propose adapting the baseline model with adapters learned for each specific task in its standard dataset, representing Phase II of our approach, as outlined in Fig. 2.
In our approach, each linear and convolutional layer is augmented with a set of LoRA, one per type of degradation, parametrized by $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ where $n$ is the index of the degradation. Following LoRA, these low-rank matrices adjust the initial frozen weights through linear combinations, as described earlier. For simplicity, we omit the subindex indicating the specific layer of the model, but each layer has a separate set of adapters. Due to the proposed pre-training approach and the subsequent specialization via adaptation, our approach achieves superior performance across diverse tasks and generalizes effectively.
@@
While LoRA adapters serve as a plug-and-play solution for pre-trained models, there is still the need to select the most suitable adapter based on the input image, when the type of degradation is unknown. To address this limitation and derive a blind all-in-one IR method, we propose to use a lightweight degradation estimator $p\left(n|x;\theta\right)$ to estimate the probability of each degradation $n$ given the input image $x$, parametrized by $\theta$. This estimator is trained on the combination of all (seen) datasets, thus learning to identify the type of degradation present in the image. As shown in Fig. 2, the probability of the estimator is used as weights to linearly combining the task-specific adapters with the baseline parameters. Specifically, given an input image $x$, the weight update for a specific layer with baseline weights $W$ and adapter weights $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ is computed as:
@@
We propose two variants of the estimator: (i) one-hot and (ii) soft-weights. The one-hot variant simply selects the adapter corresponding to the degradation with the maximum probability (equivalent to 1 with one-hot probabilities), while the soft-weight variant uses computes the weighted average as in Eq. 1.
Our approach allows the model to linearly combine the degradation-specific knowledge to address both specific and multiple distortions. Its modular and parameter efficient architecture also allows to add additional degradations with small training effort: train additional adapters for the new task, and update the estimator.
@@
Restormer [63] 14.50 .464 24.42 .781 27.12 .877
PromptIR [40] 14.28 .444 24.49 .762 27.70 .870
Ours OH 21.69 .692 27.04 .850 28.09 .907
Ours SW 19.37 .594 27.05 .850 28.09 .906

### v2

@@ Abstract
Blind all-in-one image restoration models aim to recover a high-quality image from an input degraded with unknown distortions. However, these models require all the possible degradation types to be defined during the training stage while showing limited generalization to unseen degradations, which limits their practical application in complex cases. In this paper, we introduce ABAIR, a simple yet effective adaptive blind all-in-one restoration model that not only handles multiple degradations and generalizes well to unseen distortions but also efficiently integrates new degradations by training only a small subset of parameters. We first train our baseline model on a large dataset of natural images with multiple synthetic degradations. To enhance its ability to recognize distortions, we incorporate a segmentation head that estimates per-pixel degradation types. Second, we adapt our initial model to varying image restoration tasks using independent low-rank adapters. Third, we learn to adaptively combine adapters to versatile images via a flexible and lightweight degradation estimator. This specialize-then-merge approach is both powerful in addressing specific distortions and flexible in adapting to complex tasks. Moreover, our model not only surpasses state-of-the-art performance on five- and three-task IR setups but also demonstrates superior generalization to unseen degradations and composite distortions.
@@ 1 Introduction
Figure 1: Our model significantly outperforms state-of-the-art all-in-one image restoration (IR) methods, DiffUIR [81], X-Restormer [8], and AdaIR [10], across five known IR tasks, three unseen tasks, and three mixed degradation scenarios. The plot is normalized along each axis, with the lowest value positioned on the second circle and the highest value on the outermost circle.
@@
The inherently ill-posed nature of IR presents a significant challenge for conventional approaches limiting their effectiveness [46, 27, 42, 26, 59, 13, 22, 16]. Recent advances in deep learning techniques have led to remarkable progress in IR [32, 76, 49, 51, 12], achieving substantial improvements in reconstruction accuracy. However, these frameworks demonstrate superior performance only in modeling dedicated degradations, as they are typically trained on IR datasets with specific degradations.
@@
To address the limitations of using separate models for each degradation type, recent IR models have adopted an all-in-one approach, i.e., they are designed to handle multiple degradation types within a single model, alleviating the dependency on dedicated models for various IR tasks. While these models have demonstrated success in integrating diverse IR tasks into a unified framework, several challenges persist when applied to real-world IR problems. First, these models assume that the type of degradations are specified in advance, allowing them to target and remove them specifically. In practice, however, IR must operate in a blind setting, with no prior knowledge of the degradation present in a given image, making blind IR significantly more challenging. Second, real images frequently contain composite degradations — e.g., a moving vehicle in a low-light scene, or a compressed image captured in hazy weather. Although existing models are designed to handle diverse degradation types, they typically process only a single type of distortion at a time, limiting their effectiveness in practical cases. Third, current all-in-one methods require access to all degradation types during training, which restricts their generalizability to unseen degradations. Adapting these models to new, unobserved degradation types while retaining the all-in-one functionality typically necessitates retraining the entire model with the expanded set of degradations — a process that is both computationally expensive and time-consuming.
@@
Contribution: This paper presents an Adaptive Blind All-in-One Image Restoration (ABAIR) method, designed to bridge the gap between IR techniques and their application in practical complex scenarios. ABAIR effectively addresses multiple and composite degradations with a flexible structure that can be easily updated to include new degradations. Our approach is a simple yet effective scheme that combines three main components. First, we propose a large-scale pretraining with synthetic degradations to obtain a robust weight initialization. To increase variability, we introduce Degradation CutMix [74], a technique for generating images with controlled mixtures of degradations. Additionally, we enhance this pretraining by incorporating a segmentation head to recognize per-pixel degradation types. Second, we train independent restoration adapters — specifically LoRA [23] — to bridge the gap between synthetic and real data. Third, we train a lightweight image-degradation estimator to select the most suitable combination of adapters, enabling a blind all-in-one IR method. By leveraging the synthetic pretraining and the specialize-then-merge scheme, our model achieves handling composite distortions within a single image. Moreover, our pretraining and adapter-based design provides a flexible and extensible model structure that can be easily updated to address new distortions. By training a new adapter for the added degradation and retraining the lightweight estimator, our approach preserves prior knowledge of other IR tasks without requiring full model retraining. This enables a blind all-in-one model for versatile IR, achieving state-of-the-art performance on all-in-one IR benchmarks, and generalizing to unseen and mixed degradations, as shown in Figure 1.
@@ 2 Related Work
Recently, multi-degradation and all-in-one IR approaches have gained significant attention. Multi-degradation methods [76, 4, 75, 73, 48, 58, 19, 38, 18, 9] proposed model architectures that are effective on distinct IR tasks. However, these methods are trained so that one set of parameters can handle only one specific type of degradation. Therefore, one still needs to allocate different weights for different degradations. Furthermore, to restore a versatile image, the type of degradation has to be known so that the corresponding parameters can be loaded, this non-blind scheme further hinders the efficiency and effectiveness for real-world applications.
@@
Meanwhile, blind all-in-one IR approaches utilize specialized modules to distinguish degradation types blindly. For instance, AirNet [32] and OneRestore [20] use a contrastive loss to extract a latent degradation representation from the input images. X-Restormer [8] adds a spatial self-attention module to the transformer block to enhance spatial mapping capabilities. AdaIR [10] operates in the frequency domain to distinguish the distinct degradation types while IDR [78] learns degradation-specific priors and incorporates them into restoration. Despite the ability to blindly process images, these methods are still incapable of dealing with composite distortions or unseen distortions, due to their focus on dedicated degradations and standard IR benchmarks.
@@
Figure 2: General schema of our proposed method. Our method is divided into three phases. In Phase I we pretrain our baseline with synthetic degradations on high-fidelity images. Each image contains different degradations in different regions — Degradation CutMix [74], and a segmentation head learns to predict them, while a restoration loss aims at restoring the image. In this way, the model can distinguish and generalize well to multiple degradations in a controlled manner. In Phase II, we learn degradation-specific adaptors using standard image restoration datasets. In Phase III, we learn a lightweight degradation estimator to adaptively blend the adapters based on the degradation profile of the input image. This three-phase methodology makes our method flexible to deal with images containing multiple degradations and easy to update for new ones as it only requires training an adapter for the new distortion and retraining the lightweight estimator.
@@
Prompt learning techniques, which capture task-specific context, have shown promise in guiding adaptation for vision tasks [24]. In all-in-one IR, PromptIR [49] integrates a dedicated prompt block to capture degradation-specific features from input images, while DA-CLIP [37], MPerceiver [2], ProRes [39] and Painter [62] leverage large pretrained models as prompt generators. However, the latter approaches, as well as diffusion-based methods [81, 25, 35], are often constrained by the high memory demands of large models and iterative processes, respectively. Additionally, existing all-in-one approaches require all degradation types to be predefined during training, limiting the addition of new distortion to already trained models. All-in-one IR also places great emphasis on adverse weather removal [60, 47, 84, 85], compression [77, 52, 33], video [80] or techniques to improve existing architectures [50, 66].
In contrast, aiming at simple and versatile image restoration, we propose a three-phase approach to blindly process images with various degradations, either in single or composite forms. First, we introduce a synthetic pretraining stage to establish a strong baseline model. Next, we employ a specialize-then-merge strategy, where the model is first adapted to individual real-world degradations before learning to address the problem in a blind all-in-one manner.
@@
(f) Deg. Cutmix
Figure 3: Examples of our synthetic degradation generation for five traditional degradations (a)-(e) and our Degradation Cutmix (f).
@@
Fine-tuning all model parameters is computationally expensive, especially for large-scale models, but Parameter Efficient Fine-Tuning (PEFT) addresses this by reducing the number of trainable parameters and memory usage while maintaining performance [67]. Low-Rank Adaptation (LoRA) [23] introduced a reparametrization strategy for fine-tuning, where the weights of a specific layer $W^{\prime}=W+\Delta W$ are represented as a linear combination of the pretrained weights $W$ and the updated term $\Delta W$ obtained through a low-rank decomposition, $\Delta W=BA$. Here, $B\in\mathbb{R}^{d\times k}$, $A\in\mathbb{R}^{r\times k}$ with rank $r\ll\text{min}(d,k)$. By optimizing only the low-rank matrices $A$ and $B$, LoRA reduces the number of parameters required for adapting that layer, minimizing memory and computational demands. Building upon LoRA, alternative decomposition methods have been proposed, such as VeRA [29] and Conv-LoRA[82].
PEFT techniques have proven valuable not only for domain adaptation in large models but also for applications like task arithmetic and continual learning, mitigating the problem of catastrophic forgetting — where models lose previously acquired knowledge upon learning new tasks [11]. In IR, Park et al. [48] applied low-rank decomposition within single-task IR frameworks to enhance model performance. Conversely, we propose a simple approach to merge different adaptors to obtain a flexible blind all-in-one model capable of learning new tasks without forgetting previously learned IR tasks. This is achieved through a disentangled scheme that requires minimal retraining for each new task.
@@ 3 Method
Figure 2 presents an overview of our proposed approach for enhancing low-quality images by systematically addressing the key limitations in current IR methods. Our approach consists of three phases, roughly targeting the following challenges: (i) robust generalization to varying types of degradations (Phase I), (ii) effective adaptation to specific degradations (Phase II), (iii) flexible blind all-in-one mechanism able to handle unknown and mixed degradations (Phase III). Note that we freeze the parameters of the baseline after Phase I and adapters after Phase II, preserving maximal knowledge learned from previous phases and avoiding catastrophic forgetting. This design also allows us to adapt to new distortions with minimal training as only a new adapter and the lightweight estimator will be trained.
Phase I: pretraining with synthetic degradations.
Recent advances demonstrate that large-scale pretrained models can significantly improve performance across various tasks [1]. Moreover, pretraining with synthetic data has also shown effectiveness when the domain gap is properly addressed [68]. We hypothesize that applying large-scale pretraining to IR can yield a notable performance boost compared to training solely on traditional IR datasets. Thus, as depicted in Figure 2, Phase I involves training a baseline IR model with large-scale data and synthetic degradations to provide a robust weight initialization.
To obtain large-scale data for pretraining, we define a degradation pipeline that generates a low-quality counterpart for each high-quality input image by introducing different synthetic degradations. We pre-train on five distortions: noise, blur, rain, haze, and low-light conditions, as: (i) they are the most studied and commonly observed in real-world scenarios, (ii) they provide a diverse and balanced mix of low- and high-frequency degradations, and (iii) they enable a strong foundation for other distortions. Each synthetic distortion closely approximates its real-world counterpart. For instance, low-light conditions are simulated by compressing the image histogram, reducing the dynamic range of pixel intensities. In contrast, haze is introduced by adding an achromatic layer based on the depth map estimated from DepthAnythingv2 [68]. Figure 3 (a)-(e) illustrate examples of each synthetic degradation. We use the Google Landmarks dataset (GLD) [65], a large-scale collection of five million images of diverse landmarks worldwide. To ensure that the input images are high quality, we filter out the images with a resolution smaller than 400 pixels and exclude images with a NIMA score [57] below the 4.90 threshold.
@@
While training on single distortions allows the model to handle each type individually, real-world IR often involves mixed degradations. However, applying multiple synthetic degradations simultaneously can severely deteriorate image quality, making it harder for the model to reconstruct the original image (see Table 6). To address this, we implement two strategies: (i) a Degradation CutMix technique, and (ii) a segmentation head with a cross-entropy loss. In Figure 3 (f), we show our first strategy — inspired by Yang et al. [68] for depth estimation, where two different degradations are applied in separated random regions. Although this setup is not fully realistic, it helps the model distinguish and manage multiple degradation types within the same image. Our second strategy incorporates a segmentation head into the baseline model, producing a per-pixel distortion map. By comparing this map with the ground truth map, we guide the model to recognize and differentiate multiple types of degradation. The segmentation head is removed in subsequent phases. Additional details on our baseline and synthetic degradation generation are provided in the Supplementary Material.
Table 1: 5-degradations setup. Quantitative results on five IR datasets comparing the state-of-the-art all-in-one methods and our approach.
@@
AdaIR [10] 38.02 .981 30.53 .978 31.35 .889 28.12 .858 23.00 .845 30.20 .910 28M
Ours OH 38.18 .983 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.24 .921 41M
Ours SW 38.22 .984 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.25 .921 41M
@@
While synthetic degradations provide a robust weight initialization for our baseline model, a domain shift remains between the pretraining data (due to the different dataset and synthetic degradations), standard IR datasets, and real-world conditions. For instance, accurately simulating haze is particularly challenging, as it depends on light scattering by particles, which varies with depth and atmospheric conditions. To mitigate this gap, we introduce Phase II of our approach, where the baseline model is adapted using task-specific adapters specializing on a single degradation by using a standard IR dataset, as outlined in Figure 2.
In this phase, we augment each linear and convolutional layer with a set of LoRA adapters, one per degradation type and parametrized by $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ where $n$ denotes the degradation index. These low-rank matrices refine the initial frozen weights through linear combinations and are trained individually for each degradation type. For simplicity, we omit the layer-specific indices, but each layer has its own set of adapters. By leveraging pretraining and subsequent specialization via adapters, our approach achieves superior performance across diverse tasks and exhibits improved generalization.
@@
LoRA adapters serve as a plug-and-play solution to refine our pretrained model. However, we still need to select the most suitable adapter based on the input image, when the degradation type is unknown. To address this limitation and derive a blind all-in-one IR method, we propose to use a lightweight degradation estimator $p\left(n|x;\theta\right)$ to estimate the probability of each degradation $n$ given the input image $x$, parametrized by $\theta$. This estimator is trained to identify the type of degradation in the image using the combination of all the IR datasets in Phase II. As shown in Figure 2, the probabilities of the estimator are used to linearly combine the task-specific adapters with the baseline weights. In this way, the baseline is adapted with the task-specific LoRAs according to the lightweight estimator and, therefore, each input image individually. Specifically, given an input image $x$, the weight update for a specific layer with baseline weights $W$ and adapter weights $\left\{A_{n}\right\}$ and $\left\{B_{n}\right\}$ is computed as:
@@
We propose two variants of the estimator: (i) one-hot (OH) and (ii) soft-weights (SW). The one-hot variant simply selects the adapter corresponding to the degradation with the maximum probability, while the soft-weight variant computes the weighted average as in Eq. 1.
With a strong baseline model capable of generalizing effectively, our approach not only achieves superior performance on known degradations but also performs well on unseen and mixed degradations. Moreover, thanks to its modular architecture, we empirically demonstrate that performance on unseen degradation types can be further improved with minimal training effort. Specifically, one only needs to train an additional adapter for the new degradation type and retrain the lightweight estimator — together requiring fewer than four million parameters.
@@
Restormer [76] 14.50 .464 24.42 .781 26.12 .877
AdaIR [10] 14.13 .438 24.43 .771 26.54 .867
PromptIR [49] 14.28 .444 24.49 .762 26.70 .870
Ours OH 18.76 .612 27.04 .850 28.09 .907
Ours SW 17.13 .564 27.05 .850 28.09 .906